# Learning Individual Behavior in Agent-Based Models with Graph Diffusion Networks

### v1

@@ C.1 Experimental design
In our experiments, we considered three combinations of the parameter $\xi$ for the Schelling ABM and four combinations of the matrix $\Psi$ for the Predator-Prey ABM. For each parameter setting, we generated 8 training datasets obtained with different initial seeds and trained a surrogate model and ablated model for each. In total we trained 112 models, 64 ($8\times 4\times 2$) for Predator-Prey, and 48 ($8\times 3\times 2$) for Schelling. All our evaluations are done across these 8 models per parameter configuration.
We fixed some of the ABM parameters across experiments, which are reported in Table 6. For Predator-Prey, density refers to the density of agents that are initialized as Alive, whereas the number of agents refers to the total number of agents in the system (Alive, Dead, Pregnant and Unborn agents). Color Ratio indicates the ratio between the number of black and white agents; Kind Ratio indicates the ratio between the number of predators and the number of preys.
@@
Micro evaluation. We evaluate how well the surrogate captures individual behavior of agents on a future ramification dataset of $T=25$ timesteps. We generate this out-of-training dataset by giving as initial condition the last system configuration $\mathbf{Z}_{T-1}[r=0]$ from the training ramification dataset. Thus, for each agent $i\in A$ we have 24 initial conditions $(\mathbf{Z}_{t}^{(i)},\{\mathbf{Z}_{t}^{(j)}\}_{j\in N_{t}^{(i)}})$ and 500 outcomes $\mathbf{Z}_{t+1}^{(i)}$ to build 24 ground truth probability distributions, such as equation (1). Then, we use our surrogate model to generate 500 outcomes $\mathbf{Z}_{t+1}^{(i)}$ given as condition $(\mathbf{Z}_{t}^{(i)},\{\mathbf{Z}_{t}^{(j)}\}_{j\in N_{t}^{(i)}})$, producing 24 probability distributions such as (1) for each agent $i\in A$, which we compare to the probability distributions from the ground truth. For Schelling, we measure the EMD on the marginals of the coordinates $x$ and $y$. For Predator-Prey, we measure the EMD on the distributions of the phases $f_{t}^{(i)}$, and fix the distance between the different phases to 1, since they represent a categorical variable. For each of the 8 experiments, we calculate the mean EMD over all agents and all timesteps. For Schelling, we evaluate the EMD on 1950 distributions (one per agent) for the $x$ coordinate and 1950 distributions for the $y$ coordinate, over 24 timestep, yielding 93600 EMD entries. For Predator-Prey, we evaluate the EMD on 2048 distributions (one per agent) over 24 timesteps, yielding 49152 EMD entries. The box plots in Figure 4 have as entries the 8 mean EMD values for the surrogate, and 8 mean EMD values for the ablated model.
Macro evaluation. We evaluate how well the surrogate model reproduces system-level behavior by tracking summary statistics over time. We generate 100 independent simulations of 25 timesteps beyond the training horizon for each model, and compute the symmetric mean absolute error (sMAPE) between the mean ground-truth trajectory and the mean surrogate-generated trajectory across these 100 independent simulations. In the case of Schelling, where we only track the number of number of happy agents, the sMAPE definition is straightforward. Let $A_{t}$ and $F_{t}$ be the mean number of happy agents across the 100 independent simulations from the ground-truth and the surrogate model respectively. Then, the sMAPE is computed as:
@@
For each of the 8 experiments, we calculate the sMAPE over 25 timesteps and then compute its mean. The box plots in Figure 4 have as entries the 8 mean sMAPE values for the surrogate, and 8 mean EMD values of the ablated model.
@@ C.2 Reproducing emergent segregation
Figure 2 in the main text illustrated how, for three selected time steps ($t=0$, $t=15$, and $t=30$) and three different values of the intolerance threshold $\xi$, our surrogate model successfully reproduced the qualitative dynamics of the original ABM. In contrast, the ablated model failed to capture these dynamics.
Figures 5, 6, and 7 provide a more detailed view of the system’s evolution for the values $\xi_{1}$, $\xi_{2}$, and $\xi_{3}$, respectively. In addition to the previously shown snapshots, these supplementary figures include intermediate time steps ($t=5$, $t=10$, $t=20$, and $t=25$), offering a more complete picture of the dynamics.
Figure 5: Evolution of the position of black and red agents in the Schelling model, for tolerance thresholds $\xi=\xi_{1}=0.625$. We compare the ground truth (top row) with our surrogate (middle row) and the ablation (bottom row).
Figure 5 illustrates that the original ABM rapidly evolves toward a segregated configuration, which becomes visible as early as $t=5$. However, the resulting clusters remain relatively small, indicating that a low intolerance threshold allows for a moderate level of social mixing. This motivates the label “Low Segregation” for the $\xi_{1}$ parameter setting in Figure 4. The surrogate model accurately reproduces these qualitative patterns, albeit with slightly larger clusters. In contrast, the ablation model fails to capture the emergent spatial structure, remaining close to the initial random configuration across time steps.
Figure 6: Evolution of the position of black and red agents in the Schelling model, for tolerance thresholds $\xi=\xi_{2}=0.75$. We compare the ground truth (top row) with our surrogate (middle row) and the ablation (bottom row).
Figure 6 shows that, for $\xi=\xi_{2}$, the original ABM also converges toward a segregated state, but the convergence occurs more slowly than in the $\xi_{1}$ case. The resulting clusters are larger and more distinct, reflecting a “High Segregation” scenario. The surrogate model again successfully replicates these dynamics, while the ablation model continues to exhibit no structured behavior.
Figure 7: Evolution of the position of black and red agents in the Schelling model, for tolerance thresholds $\xi=\xi_{3}=0.875$. We compare the ground truth (top row) with our surrogate (middle row) and the ablation (bottom row).
Finally, Figure 7 demonstrates that, for $\xi=\xi_{3}$, the original ABM does not converge to a stable configuration. Instead, the high intolerance threshold causes agents to continuously relocate, preventing the emergence of segregated clusters. In this degenerate case, both the surrogate and the ablation models correctly reproduce the persistent disorder of the original dynamics.
@@ C.3 Reproducing emergent oscillations in predator-prey dynamics
Figure 8: Forecasting macro-level summary statistics (here, the number of alive preys and predators over time), starting from the last condition seen in training, for 100 independent simulation runs, under configuration $\Psi_{2}$ (monotonic dynamics for both predators and preys, top) and $\Psi_{3}$ (oscillations only for predators, bottom). Left: original ABM simulations. Center: surrogate. Right: ablation. The dashed vertical line indicates the end of the training phase for the surrogate and ablation models.
@@
Figure 8 extends the previous analysis to parameter sets $\Psi_{2}$ and $\Psi_{3}$, which produce, respectively, a monotonic decline in prey and predator populations, and oscillatory dynamics for preys but not for predators. The case of $\Psi_{2}$ is particularly illustrative: despite its apparent simplicity, the ablated model fails to reproduce the monotonic trends, showing an unrealistic spike in both populations immediately after the training window. In contrast, the surrogate model accurately captures the expected decay. Under $\Psi_{3}$, the surrogate successfully replicates the oscillations in the prey population and the stable predator trend, while the ablation once again outputs generic dynamics, largely insensitive to the underlying regime. This highlights the ablation model’s inability to distinguish between qualitatively different behaviors.
@@
Beyond aggregate population counts, it is instructive to examine the spatial distribution of predators and preys over time. Prior studies [8, 9] have shown that similar predator-prey models give rise to rich spatial dynamics, characterized by the spontaneous emergence of structured patterns (see, e.g., Figure 1.3 in [16]). Starting from initially random configurations, the interactions between agents give rise to both short- and long-range spatial correlations, with predators and preys organizing into dynamic clusters and propagating waves. These patterns are reminiscent of those observed in spatially extended reaction-diffusion systems and bear a strong resemblance to spatial chaos phenomena in evolutionary game theory [28].
@@
We observe similar spatial patterns in our predator-prey ABM. Figures 9, 10, 11, and 12 show the evolution of the positions of predators and preys on the grid.
@@
Figure 9: Evolution of the position of preys (black) and predators (red) in the predator-prey ABM, for parameters $\Psi=\Psi_{1}$. We compare the ground truth (top row) with our surrogate (middle row) and the ablation (bottom row).
@@
Figure 9 illustrates the spatial dynamics under parameter set $\Psi=\Psi_{1}$, which is characterized by oscillations in predator and prey populations. The ground-truth model displays a rise in population densities mid-simulation, followed by a near-extinction phase toward the end, in line with the temporal trends shown in Figure 3. Notably, the ground-truth also exhibits distinct spatial patterns, with predators and preys forming dynamic clusters. The surrogate model successfully reproduces both the population dynamics and the emergent spatial structures, whereas the ablated model fails to capture any meaningful spatial organization.
@@
Figure 10: Evolution of the position of preys (black) and predators (red) in the predator-prey ABM, for parameters $\Psi=\Psi_{2}$. We compare the ground truth (top row) with our surrogate (middle row) and the ablation (bottom row).
@@
Figure 11: Evolution of the position of preys (black) and predators (red) in the predator-prey ABM, for parameters $\Psi=\Psi_{3}$. We compare the ground truth (top row) with our surrogate (middle row) and the ablation (bottom row).
@@
Figure 12: Evolution of the position of preys (black) and predators (red) in the predator-prey ABM, for parameters $\Psi=\Psi_{4}$. We compare the ground truth (top row) with our surrogate (middle row) and the ablation (bottom row).
@@
Similar results are observed in Figures 10, 11, and 12, which capture different dynamic regimes for predators and preys. In all cases, the surrogate model accurately reproduces both spatial and temporal patterns, while the ablated model consistently fails to capture the underlying dynamics or spatial structure.
@@ C.4 Quantitative results
Figure 13: Example distributions of the position of a single agent in the Schelling model, in two different conditions: (a) happy agent; (b) unhappy agent. Histograms on the left represents the original ground truth ABM, those on the right the ones obtained by our surrogate model. The dashed red vertical line indicates the initial position of the agent. Note that the coordinates have been rescaled to [-25, 25], compared to the [0, 50] interval used for easiness of exposition in Algorithm 2.
In this section, we present additional details regarding the quantitative evaluation discussed in Section 4.2 (page 4.2). In particular here we detail further the micro-level evaluation; that is, how much our surrogate model is able to reproduce the distribution of possible states of an agent at $t$ (that is, $\mathbf{Z}_{t}^{{}^{(i)}}$) given its past state $\mathbf{Z}_{t-1}^{{}^{(i)}}$. To better understand how this comparison works, Figure 13 (left side) shows the distribution for a single agent of the position $x_{t}^{(i)}$, one of the components of the state $\mathbf{Z}_{t}^{{}^{(i)}}$, for each of the two possible past conditions in the Schelling model, happy and unhappy. We see that when the agent is happy, its position remains fixed, while in the case the agent is unhappy, it moves randomly with a certain distribution peaked around the starting point (implicitly defined by the ABM, see Algorithm 2). Our surrogate model aims at reproducing this distribution, without knowing the original ABM and without access to the latent variable happy/unhappy, but only observing the sequence of states and the graph of interactions. The distributions obtained by our surrogate in the same starting conditions are shown on the right side of the figure. To evaluate the quality of this reconstruction, we quantify it as the earth mover distance between the distributions, for each agent in each timestep, and then aggregate these measures.
@@
Schelling model. Figure 14 shows the distribution of the EMD scores obtained by our surrogate model and by the ablation model only for the stochastic rules of the Schelling model. In this ABM, the stochasticity lies in the random movement of the unhappy agents (Algorithm 2, lines 16-25). We see that our surrogate is able to capture these distributions well (EMD averages below 5, considering that the scale is given by the length of the grid, $L=50$) and better than the surrogate model.
@@
Predator-prey model. Figure 15 instead shows the distribution of the EMD scores only for the stochastic rules of the Predator-Prey model. Here, the stochasticity lies in the Alive phase, where the agent might transition to another phase depending on its interaction with its neighbors. We see that also in this model our surrogate is able to capture the stochasticity well, with EMD averaging below 0.1 for each of the four configurations $\Psi_{1-4}$. Here the scale is in $[0,1]$ since this error is measured on the life phase binary vector. We further investigate these results in Figure 16 by dividing them by agent state in each configuration. Here the stochastic transitions happen only in the first column (Alive). This figure confirms that the EMD values are quite low, and lower than those obtained by the ablation model. Moreover, here we observe that also the deterministic transitions are recovered with precision (error is always below $10^{-3}$)) by our surrogate model.
@@
Figure 14: Distribution of errors (EMD) obtained by our surrogate model (top row) and by the ablation model (bottom row) only for the stochastic rules of the Schelling ABM for each considered configuration of the parameter $\xi$ of the ABM. Different colors indicate independent experiments that only differ by the random seed used to generate the ground truth data.
@@
Figure 15: Distribution of errors (EMD) obtained by our surrogate model (top row) and by the ablation model (bottom row) only for the stochastic rules of the Predator-Prey ABM for each considered configuration of the parameter matrix $\Psi$ of the ABM. Different colors indicate independent experiments that only differ by the random seed used to generate the ground truth data.
@@
Figure 16: Mean error (EMD) obtained by our surrogate (on the left) and by the ablation model (on the right) for each configuration of the Predator-Prey ABM (on the rows, $\Psi_{1-4}$) and for each initial state of an agent (on the columns).

### v2

@@ C.1 Experimental design
In our experiments, we considered three combinations of the parameter $\xi$ for the Schelling ABM and four combinations of the matrix $\Psi$ for the Predator-Prey ABM. For each parameter setting, we generated 8 training datasets obtained with different initial seeds and trained a surrogate model and two ablated models for each. In total, we trained 168 models, 96 ($8\times 4\times 3$) for Predator-Prey, and 72 ($8\times 3\times 3$) for Schelling. All our evaluations are done across these 8 models per parameter configuration.
We fixed some of the ABM parameters across experiments, which are reported in Table 9. For Predator-Prey, density refers to the density of agents that are initialized as Alive, whereas the number of agents refers to the total number of agents in the system (Alive, Dead, Pregnant, and Unborn agents). Color Ratio indicates the ratio between the number of black and white agents; Kind Ratio indicates the ratio between the number of predators and the number of preys.
@@
Micro evaluation. We evaluate how well the surrogate captures individual behavior of agents on a future ramification dataset of $T=25$ timesteps. We generate this out-of-training dataset by giving as initial condition the last system configuration $\mathbf{Z}_{T-1}[r=0]$ from the training ramification dataset. Thus, for each agent $i\in A$ we have 24 initial conditions $(\mathbf{Z}_{t}^{(i)},\{\mathbf{Z}_{t}^{(j)}\}_{j\in N_{t}^{(i)}})$ and 500 outcomes $\mathbf{Z}_{t+1}^{(i)}$ to build 24 ground truth probability distributions, such as equation (1). Then, we use our surrogate model to generate 500 outcomes $\mathbf{Z}_{t+1}^{(i)}$ given as condition $(\mathbf{Z}_{t}^{(i)},\{\mathbf{Z}_{t}^{(j)}\}_{j\in N_{t}^{(i)}})$, producing 24 probability distributions such as (1) for each agent $i\in A$, which we compare to the probability distributions from the ground truth. For Schelling, we measure the EMD on the marginals of the coordinates $x$ and $y$. For Predator-Prey, we measure the EMD on the distributions of the phases $f_{t}^{(i)}$, and fix the distance between the different phases to 1, since they represent a categorical variable. For each of the 8 experiments, we calculate the mean EMD over all agents and all timesteps. For Schelling, we evaluate the EMD on 1950 distributions (one per agent) for the $x$ coordinate and 1950 distributions for the $y$ coordinate, over 24 timesteps, yielding 93600 EMD entries. For Predator-Prey, we evaluate the EMD on 2048 distributions (one per agent) over 24 timesteps, yielding 49152 EMD entries. The box plots in Figure 4 have as entries the 8 mean EMD values for the surrogate, 8 mean EMD values for the diffusion-only ablated model, and 8 mean EMD values for the GNN-only ablated model.
Macro evaluation. We evaluate how well the surrogate model reproduces system-level behavior by tracking summary statistics over time. We generate 100 independent simulations of 25 timesteps beyond the training horizon for each model, and compute the symmetric mean absolute error (sMAPE) between the mean ground-truth trajectory and the mean surrogate-generated trajectory across these 100 independent simulations. In the case of Schelling, where we only track the number of happy agents, the sMAPE definition is straightforward. Let $A_{t}$ and $F_{t}$ be the mean number of happy agents across the 100 independent simulations from the ground-truth and the surrogate model respectively. Then, the sMAPE is computed as:
@@
For each of the 8 experiments, we calculate the sMAPE over 25 timesteps and then compute its mean. The box plots in Figure 4 have as entries the 8 mean sMAPE values for the surrogate, 8 mean sMAPE values for the diffusion-only ablated model, and 8 mean sMAPE values for the GNN-only ablated model.
@@ C.2 Reproducing emergent segregation
Figure 2 in the main text illustrated how, for three selected time steps ($t=0$, $t=15$, and $t=30$) and three different values of the intolerance threshold $\xi$, our surrogate model successfully reproduced the qualitative dynamics of the original ABM. In contrast, the ablated models failed to capture these dynamics.
Figures 8, 9, and 10 provide a more detailed view of the system’s evolution for the values $\xi_{1}$, $\xi_{2}$, and $\xi_{3}$, respectively. In addition to the previously shown snapshots, these supplementary figures include intermediate time steps ($t=5$, $t=10$, $t=20$, and $t=25$), offering a more complete picture of the dynamics.
Figure 8: Evolution of the position of black and red agents in the Schelling model, for tolerance thresholds $\xi=\xi_{1}=0.625$. We compare the ground truth (top row) with our surrogate (second row) and the ablations (third and fourth row).
Figure 8 illustrates that the original ABM rapidly evolves toward a segregated configuration, which becomes visible as early as $t=5$. However, the resulting clusters remain relatively small, indicating that a low intolerance threshold allows for a moderate level of social mixing. This motivates the label “Low Segregation” for the $\xi_{1}$ parameter setting in Figure 4. The surrogate model accurately reproduces these qualitative patterns, albeit with slightly larger clusters. In contrast, the diffusion-only ablation fails to capture the emergent spatial structure, remaining close to the initial random configuration across time steps. Instead, the GNN-only ablation ends up overlapping most agents on the same coordinates.
Figure 9: Evolution of the position of black and red agents in the Schelling model, for tolerance thresholds $\xi=\xi_{2}=0.75$. We compare the ground truth (top row) with our surrogate (second row) and the ablations (third and fourth row).
Figure 9 shows that, for $\xi=\xi_{2}$, the original ABM also converges toward a segregated state, but the convergence occurs more slowly than in the $\xi_{1}$ case. The resulting clusters are larger and more distinct, reflecting a “High Segregation” scenario. The surrogate model again successfully replicates these dynamics, while the ablation models continue to exhibit no structured behavior.
Figure 10: Evolution of the position of black and red agents in the Schelling model, for tolerance thresholds $\xi=\xi_{3}=0.875$. We compare the ground truth (top row) with our surrogate (second row) and the ablations (third and fourth row).
Finally, Figure 10 demonstrates that, for $\xi=\xi_{3}$, the original ABM does not converge to a stable configuration. Instead, the high intolerance threshold causes agents to continuously relocate, preventing the emergence of segregated clusters. In this degenerate case, both the surrogate and the diffusion-only ablation model correctly reproduce the persistent disorder of the original dynamics, while the GNN-only ablation still fails.
@@ C.3 Reproducing emergent oscillations in predator-prey dynamics
Figure 11: Forecasting macro-level summary statistics (here, the number of alive preys and predators over time), starting from the last condition seen in training, for 100 independent simulation runs, under configuration $\Psi_{2}$ (monotonic dynamics for both predators and preys, top) and $\Psi_{3}$ (oscillations only for predators, bottom). First colummn: original ABM simulations. Second column: surrogate. Third and fourth column: ablations. The dashed vertical line indicates the end of the training phase for the surrogate and ablation models.
@@
Figure 11 extends the previous analysis to parameter sets $\Psi_{2}$ and $\Psi_{3}$, which produce, respectively, a monotonic decline in prey and predator populations, and oscillatory dynamics for preys but not for predators. The case of $\Psi_{2}$ is particularly illustrative: despite its apparent simplicity, the ablated models fail to reproduce the monotonic trends, showing an unrealistic spike in both populations immediately after the training window in the diffusion-only ablated model, and a roughly constant number of preys and predators in the GNN-only ablated model. In contrast, the surrogate model accurately captures the expected decay. Under $\Psi_{3}$, the surrogate successfully replicates the oscillations in the prey population and the stable predator trend, while the diffusion-only ablation once again outputs generic dynamics, largely insensitive to the underlying regime, and the GNN-only generates an oscillating dynamic for both. This highlights the ablation models’ inability to distinguish between qualitatively different behaviors.
@@
Beyond aggregate population counts, it is instructive to examine the spatial distribution of predators and preys over time. Prior studies [1, 2] have shown that similar predator-prey models give rise to rich spatial dynamics, characterized by the spontaneous emergence of structured patterns (see, e.g., Figure 1.3 in [3]). Starting from initially random configurations, the interactions between agents give rise to both short- and long-range spatial correlations, with predators and preys organizing into dynamic clusters and propagating waves. These patterns are reminiscent of those observed in spatially extended reaction-diffusion systems and bear a strong resemblance to spatial chaos phenomena in evolutionary game theory [nowak1992evolutionary].
@@
We observe similar spatial patterns in our predator-prey ABM. Figures 12, 13, 14, and 15 show the evolution of the positions of predators and preys on the grid.
@@
Figure 12: Evolution of the position of preys (black) and predators (red) in the predator-prey ABM, for parameters $\Psi=\Psi_{1}$. We compare the ground truth (top row) with our surrogate (second row) and the ablations (third and fourth row).
@@
Figure 12 illustrates the spatial dynamics under parameter set $\Psi=\Psi_{1}$, which is characterized by oscillations in predator and prey populations. The ground-truth model displays a rise in population densities mid-simulation, followed by a near-extinction phase toward the end, in line with the temporal trends shown in Figure 3. Notably, the ground-truth also exhibits distinct spatial patterns, with predators and preys forming dynamic clusters. The surrogate model successfully reproduces both the population dynamics and the emergent spatial structures, whereas the ablated models fail to capture any meaningful spatial organization.
@@
Figure 13: Evolution of the position of preys (black) and predators (red) in the predator-prey ABM, for parameters $\Psi=\Psi_{2}$. We compare the ground truth (top row) with our surrogate (second row) and the ablations (third and fourth row).
@@
Figure 14: Evolution of the position of preys (black) and predators (red) in the predator-prey ABM, for parameters $\Psi=\Psi_{3}$. We compare the ground truth (top row) with our surrogate (second row) and the ablations (third and fourth row).
@@
Figure 15: Evolution of the position of preys (black) and predators (red) in the predator-prey ABM, for parameters $\Psi=\Psi_{4}$. We compare the ground truth (top row) with our surrogate (second row) and the ablations (third and fourth row).
@@
Similar results are observed in Figures 13, 14, and 15, which capture different dynamic regimes for predators and preys. In all cases, the surrogate model accurately reproduces both spatial and temporal patterns, while the ablated models consistently fail to capture the underlying dynamics or spatial structure.
@@ C.4 Quantitative results
Figure 16: Example distributions of the position of a single agent in the Schelling model, in two different conditions: (a) happy agent; (b) unhappy agent. Histograms on the left represent the original ground-truth ABM, those on the right the ones obtained by our surrogate model. The dashed red vertical line indicates the initial position of the agent. Note that the coordinates have been rescaled to [-25, 25], compared to the [0, 50] interval used for ease of exposition in Algorithm 2.
In this section, we present additional details regarding the quantitative evaluation discussed in Section 4.2 (page 4.2). In particular here we detail further the micro-level evaluation; that is, how much our surrogate model can reproduce the distribution of possible states of an agent at $t$ (that is, $\mathbf{Z}_{t}^{{}^{(i)}}$) given its past state $\mathbf{Z}_{t-1}^{{}^{(i)}}$. To better understand how this comparison works, Figure 16 (left side) shows the distribution for a single agent of the position $x_{t}^{(i)}$, one of the components of the state $\mathbf{Z}_{t}^{{}^{(i)}}$, for each of the two possible past conditions in the Schelling model, happy and unhappy. We see that when the agent is happy, its position remains fixed, while in the case the agent is unhappy, it moves randomly with a certain distribution peaked around the starting point (implicitly defined by the ABM, see Algorithm 2). Our surrogate model aims at reproducing this distribution, without knowing the original ABM and without access to the latent variable happy/unhappy, but only observing the sequence of states and the graph of interactions. The distributions obtained by our surrogate in the same starting conditions are shown on the right side of the figure. To evaluate the quality of this reconstruction, we quantify it as the Earth Mover’s Distance between the distributions, for each agent in each timestep, and then aggregate these measures.
@@
Schelling model. Figure 17 shows the distribution of the EMD scores obtained by our surrogate model and by the ablation model only for the stochastic rules of the Schelling model. In this ABM, the stochasticity lies in the random movement of the unhappy agents (Algorithm 2, lines 16-25). We see that our surrogate can capture these distributions well (EMD averages below 5, considering that the scale is given by the length of the grid, $L=50$) and better than the surrogate model.
@@
Predator-prey model. Figure 18 instead shows the distribution of the EMD scores only for the stochastic rules of the Predator-Prey model. Here, the stochasticity lies in the Alive phase, where the agent might transition to another phase depending on its interaction with its neighbors. We see that also in this model our surrogate is able to capture the stochasticity well, with EMD averaging below 0.1 for each of the four configurations $\Psi_{1-4}$. By contrast, the diffusion-only ablation model’s EMD for these same stochastic rules averages above 0.12, up to values around 0.2. Here, the scale is in $[0,1]$ since this error is measured on the life phase binary vector. We further investigate these results in Figure 19 by dividing them by agent state in each configuration. Here, the stochastic transitions happen only in the first column (Alive). This figure confirms that the EMD values are quite low, and lower than those obtained by the ablation model. Moreover, here we observe that also the deterministic transitions are recovered with precision (error is always below $10^{-3}$)) by our surrogate model.
@@
Figure 17: Distribution of errors (EMD) obtained by our surrogate model (top row) and by the diffusion-only ablation model (bottom row) only for the stochastic rules of the Schelling ABM for each considered configuration of the parameter $\xi$ of the ABM. Different colors indicate independent experiments that only differ by the random seed used to generate the ground truth data.
@@
Figure 18: Distribution of errors (EMD) obtained by our surrogate model (top row) and by the diffusion-only ablation model (bottom row) only for the stochastic rules of the Predator-Prey ABM for each considered configuration of the parameter matrix $\Psi$ of the ABM. Different shades indicate independent experiments that only differ by the random seed used to generate the ground truth data.
@@
Figure 19: Mean error (EMD) obtained by our surrogate (on the left) and by the diffusion-only ablation model (on the right) for each configuration of the Predator-Prey ABM (on the rows, $\Psi_{1-4}$) and for each initial state of an agent (on the columns).
C.5 Macro baseline (AR1)
In this subsection, we show that a simple yet commonly used time-series baseline — the autoregressive model of order one (AR(1)) — completely fails to reproduce the macro-level dynamics of our system. We focus on the predator–prey model, although qualitatively similar results hold for the Schelling model and for higher-order time-series models (e.g., AR(2)).
The AR(1) model assumes that the value of a time series at step $t$ depends linearly on its previous value:
$x_{t}=\phi x_{t-1}+\epsilon_{t},$
where $\phi$ is the autoregressive coefficient and $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2})$ is Gaussian noise. We fit the model separately to each macro variable using least squares estimation on the training portion of the simulated data. The fitted coefficients $\hat{\phi}$ and $\hat{\sigma}$ are then used to generate out-of-sample forecasts, producing trajectories that can be directly compared with those generated by our surrogate model.
As shown in Figure 20, the AR(1) baseline simply extrapolates the last observed trend in the training data, completely failing to reproduce the non-monotonic, oscillatory patterns characteristic of the underlying dynamics — and even performing poorly in capturing monotonic trends. This limitation arises because the system’s behavior is defined at the micro level, while macro-level variables are only aggregate summaries of those micro interactions. Consequently, a surrogate model that learns from micro-level states, as ours does, is inherently better positioned to capture both the micro and the emergent macro dynamics.
(a) $\Psi_{1}$
(b) $\Psi_{2}$
(c) $\Psi_{3}$
(d) $\Psi_{4}$
Figure 20: Comparison of $\Psi_{1}$$\Psi_{4}$ under the surrogate GDN model and an AR(1) baseline.
SM References
  • [1] André M De Roos, Edward McCauley, and William G Wilson. Mobility versus density-limited predator-prey dynamics on different spatial scales. Proceedings of the Royal Society of London. Series B: Biological Sciences, 246(1316):117–122, 1991.
  • [2] Douglas D Donalson and Roger M Nisbet. Population dynamics and spatial scale: effects of system size on population persistence. Ecology, 80(8):2492–2507, 1999.
  • [3] Volker Grimm and Steven F Railsback. Individual-based modeling and ecology. In Individual-based modeling and ecology. Princeton university press, 2013.
  • [4] Alexander Quinn Nichol and Prafulla Dhariwal. Improved denoising diffusion probabilistic models. In Proceedings of the 38th International Conference on Machine Learning, volume 139, pages 8162–8171. PMLR, 2021.